# Minimal Interaction Edge Tuning: A New Paradigm for Visual Adaptation

### v2

@@ Abstract
The rapid scaling of large vision pretrained models makes fine-tuning tasks more and more difficult on edge devices with low computational resources. We explore a new visual adaptation paradigm called edge tuning, which treats large pretrained models as standalone feature extractors that run on powerful cloud servers. The fine-tuning carries out on edge devices with small networks which require low computational resources. Existing methods that are potentially suitable for our edge tuning paradigm are discussed. But, three major drawbacks hinder their application in edge tuning: low adaptation capability, large adapter network, and high information transfer overhead. To address these issues, we propose Minimal Interaction Edge Tuning, or MIET, which reveals that the sum of intermediate features from pretrained models not only has minimal information transfer but also has high adaptation capability. With a lightweight attention-based adaptor network, MIET achieves information transfer efficiency, parameter efficiency, computational and memory efficiency, and at the same time demonstrates competitive results on various visual adaptation benchmarks.
11footnotetext: J. Wu is the corresponding author.
@@ 1 Introduction
Pretraining a large model [10, 9, 3, 22] on general data and fine-tune it on downstream tasks is emerging as the mainstream methodology for computer vision tasks. Knowledge or representation learned from the pretraining makes adaptation to a target domain much easier than training from scratch.
@@
However, some paradoxes begin to emerge within this process or learning paradigm. One natural issue is: if the downstream dataset is small, we cannot afford to and should not store one new fine-tuned model for each downstream visual task. Parameter-efficient fine-tuning (PEFT) methods [14, 32, 15, 28, 19] are proposed to fine-tune large pretrained models with only a small set of trainable parameters, by either inserting small learnable blocks, adding few learnable tokens, or learning low rank decomposition of extra parameters. These methods achieve competitive results on various downstream adaptation tasks with less than 1% trainable parameters compared to full fine-tuning.
@@
While PEFT methods partially solved the storage cost of fine-tuned models, other problems remain. As the scale of pretrained models grows with a blazing fast speed, fine-tuning or even inference itself is becoming unaffordable, especially on devices with limited computational resources. On the other hand, the need for fine-tuning and inference on edge devices (e.g., personal computers, mobile phones or even cameras) rather than on the cloud AI infrastructure is increasing rapidly.
In this paper, we propose a new visual adaptation paradigm: edge tuning, which enables edge devices to fine-tune downstream visual tasks with a small edge network. The edge network merely receives intermediate features/activations from a pretrained model, such that the huge pretrained network is not required to be (either forward or backward) computed on edge devices. Upon the proposed paradigm, we design a new visual adaptation method: Minimal Interaction Edge Tuning, which is based on the side-tuning idea [27, 29]. Methods in this style potentially enables edge-tuning, but the interaction (i.e., features transferred) from the pretrained network to the edge device is too much for an edge device to handle. Our contributions can be summarized as follows:
  • We propose a new visual adaptation paradigm: edge tuning, where the pretrained model works as a service in a cloud server, and edge devices fine-tune small edge networks to transfer the pretrained large model towards specific downstream tasks.
  • To reduce the interaction between server and edge devices, we reveal that the sum of intermediate features effectively capture essence of the pretrained network, which leads to both minimal interaction and high accuracy.
  • We propose a new visual adaptation method: MIET, as an effective means for edge tuning. MIET is parameter-efficient, GPU memory-efficient, computation efficient, and enjoys a low communication overhead between a cloud server and edge devices. Experimental results on various datasets show that MIET is comparable with state-of-the art PEFT methods in visual adaptation accuracy, but enjoys significantly smaller resource consumptions.
2 New Pipeline: Edge Tuning
@@
Figure 1: Illustration of Edge Tuning. Visual samples for different tasks ($x_{1},x_{2},x_{3},\dots$) are collected and transferred to the cloud server. Within the cloud server, a large pretrained model $B$ acts as a standalone feature extractor, producing intermediate feature sets $(B,x_{i})$. A gather function compresses this set to $\mathrm{Gather}(B,x_{i})$ to achieve minimal interaction but still keeps essential information in $(B,x_{i})$ for downstream task learning. The gathered features for task $i$ is sent to an edge device $E_{i}$, where the fine-tuning is performed solely on $E_{i}$ with input $\mathrm{Gather}(B,x_{i})$. During inference, the pretrained model extracts features, gather them to a minimal level, transfers only a small chunk of bytes to the edge device, which then makes decisions with small storage, computational, and memory costs.
@@
As illustrated and described in Figure 1, the idea and structure of edge tuning are pretty simple. Computations are split between a pretrained model $B$ (residing in a cloud server) and edge devices. The pretrained model $B$ is general-purpose and pretrained with abundant data and computing resources, which is too big to infer on the edge device. The edge device is mostly likely not powerful enough to store the model $B$ itself, nor does it has the compute or GPU memory resources to forward compute $B$. The solution is to let $B$ extract features and transfer them to the edge device.
@@
Very important but less discussed, the edge device often lacks communication resources. That is, in real world it is often difficult or even impossible to transfer large chunks of data between the edge and the server, e.g., due to bandwidth or energy constraints. Given an input image $x$, $B$ extracts features $(B,x)$ and sends it to the edge device. For example, $(B,x)$ can be the set of activations of all intermediate Transformer blocks’ activations, which contains tens of times more bytes to transfer than that in the input $x$ itself. In such case communication is too heavy a burden for the edge device.
@@
Practical applications, however, are often keen to fine-tune and infer with a model for downstream tasks solely on the edge device with limited storage, compute, memory, and communication capabilities.
@@ 2.1 Recipes and Obstacles
Methods that allow to split computations between a large backbone and a small side-network has appeared in the community [27, 29], where the side-network receives features from the backbone but does not send its features back to the computation graph of the backbone. Thus, during fine-tuning, the backbone does not need to compute gradients. This split greatly reduces the compute and (GPU) memory costs of fine-tuning.
An even simpler recipe is linear evaluation, in which final features of the pretrained model $B$ is coupled with a simple fully connected (FC) layer. Fine-tuning is then learning the FC layer, where the transferred features are tiny, the storage, compute and memory costs for edge devices are small, too. Partial-$k$ tuning unfreezes the last $k$ layers of $B$ during fine-tuning. When $k$ is small (e.g., $k=1$), it might be able to fit into edge devices.
However, these methods do not yet fit the purpose of edge tuning we propose in Figure 1. The set of 5 goals in our edge tuning (small storage/parameters, small compute small memory footprint, small communication, high accuracy) cannot be met by existing methods. For example,
@@
  • Side-tuning methods [27] produces and transfers large features $(B,x)$, and the side-network may be too big for the edge device.
2.2 Minimal Interaction Edge Tuning
Our solution, as depicted in Figure 1, meets these requirements by proposing a gather function $\mathrm{Gather}$, which compresses the information in $(B,x)$ to $\mathrm{Gather}(B,x)$. The gathered output is succinct, hence achieving minimal interaction or communication. So long as $\mathrm{Gather}(B,x)$ distills essential information for the downstream task, the small network on the edge device, $E(\mathrm{Gather}(B,x))$ will achieve high accuracy. Because in Figure 1, $E_{i}$ is very small and it does not transfer any features back to $B$, the fine-tuning and inference processes resemble side-tuning methods, which are storage (parameter), compute, and memory efficient, too.
One additional benefit of our pipeline is that the pretrained model is agnostic to downstream tasks. The pretrained model $B$ does not need to store any task-specific parameters or gradients. It simply extract features for any input images. If an input image $x$ is useful in many downstream tasks (edge devices), the pretrained model $B$ only needs to compute its features $(B,x)$ once.
@@ 3 Method
where $z_{0}$ is embedding of $x$, and $z_{i}$ is the activations/features of the $i$-th Transformer block in $B$, as illustrated in Figure 2. The ‘Stack’ operator stacks or concatenates all tensors $z_{i}$ together. For example, when the pretrained network $B$ is ViT-B, one $3\times 224\times 224$ input image (0.60MB in size) is required to transfer more than 7.5MB to edge devices!
Because side tuning methods [27, 29] have achieved state-of-the-art accuracy in visual adaptation tasks, it is a natural hypothesis that $\mathrm{Stack}(z_{0},z_{1},\cdots,z_{N})$ contains enough information for downstream tasks. And, we aim to find a gather function $\mathrm{Gather^{*}}$ that can most effectively reduce its size but keep the essential information for downstream tasks.
3.1 Gather Function: Sum is All You Need
@@
Then the input for the $i+1$-th side block is $o_{i}+z_{i}$, or equivalently,
@@
To fit into our edge tuning pipeline, $B$ can only send external feature terms to edge devices, a requirement only the third term $\Sigma_{l=0}^{i}z_{l}$ satisfies. This term is developed through repeatedly applying the skip connections.
Furthermore, because side network computations ($F_{i}$) are low rank transformation to achieve parameter efficiency [29], a natural conjecture is that this third term (which is determined by the complex model $B$) contains more complicated features compared to the first two terms (which are determined by the simple computations $F_{i}$). In other words, it is reasonable to guess that the third term, $\Sigma_{l=0}^{i}z_{l}$, contains essential information in $in_{i+1}$. Please note that this conjecture is made with respect to any $i$ ($1\leq i\leq N$), not only the final block (i.e., $i=N$).
@@
This gather function leads to only $1/(N+1)$ communication overhead compared to the one in Eq. 1. Smaller input to the edge devices also means that the storage, compute, and memory costs will be reduced accordingly. As our experimental results will show in later sections, $\mathrm{Gather^{*}}(B,x)$ does contains essential information for downstream tasks, which leads to high accuracy for them.
@@ 3.2 Gather Function: Migrating to Downstream Tasks
One difficulty in handling different downstream tasks are the domain gaps. In some simple tasks with only low-level image patterns, high-level features extracted by a pretrained model may be harmful to the adaptation. In such cases, gathering all the intermediate features altogether is suboptimal.
So we migrate the extracted features to specific downstream tasks by only gathering features from the first $k$ layers, which leads to the final form of our minimal interaction gather function:
@@
This final form introduces a hyperparameter, $k$, which needs to be set before edge tuning. We select the best $k$ from a candidate set by using the validation set. The candidate set is formed by $N$, $43N/$ 4 and $N/2$. For example, for ViT-B with 12 layers, we pick $k$ from the set $[6,9,12]$.
Figure 3: Illustration of MIET and other fine-tuning methods. Adapter fine-tunes a model by inserting trainable adapter blocks. LoRA adds trainable low-rank decomposed matrices to parameters. Side-tuning trains a decoupled side network by passing intermediate features to the side network block-wise. MIET adds intermediate features together, and pass the sum to a trainable edge network to adapt to specific task.
3.3 The Proposed MIET Method
With $\mathrm{Gather^{*}}$, our Minimal Interaction Edge Tuning (MIET) method becomes clear and easy to describe. It contains a pretrained base vision transformer network $B$ and a low-rank attention edge network $\mathrm{LAE}$ (which will be introduced latter).
$Z_{\mathrm{mix}}=\mathrm{Gather^{*}}(B,x,k)\,.$ (7)
$Z_{\mathrm{out}}=\mathrm{LAE}(Z_{\mathrm{mix}})\,.$ (8)
From the result in [29], attention block with low-dimensional QKV projection is an effective and efficient structure for visual adaptation. Our experimental results also show that it is very effective for edge tuning. So we adopt this structure for edge network.
As shown in Figure 3, $\mathrm{LAE}$ contains $L$ blocks. Each block contains a layer norm [1] and a low-rank attention, where low-rank attention refers to self-attention with an extremely low QKV hidden dimension $r$ (e.g., hidden dimension equals 16 for feature dimension equals 768). Skip connection is also applied in each block. In practice, we take $L=4$ and $r=32$ for all downstream tasks, and find that such a small edge network scale is already enough for most visual adaptation tasks.
4 Experiments
Table 1: Results on VTAB-1K. “Params”: number of trainable parameters. “MACs Cloud”, “MACs Edge”: multiply–accumulate operations for cloud and edge model during fine-tuning. “GPU Mem Cloud”, “GPU Mem Edge”: GPU memory footprint within cloud and edge devices, respectively during fine-tuning with batch size 32. “Nat.”, “Spe.”, “Str.”: average accuracy for “Natural”, “Specialized”, “Structured” task groups. “Avg.”: group-wise average accuracy. Best results are in boldface.
@@ 4 Experiments
MIET (ours) 0.577 0.38 17.57 00.09 0.57 0.45 81.4 86.4 62.4 76.7

### v3

# Minimal Interaction Separated Tuning: A New Paradigm for Visual Adaptation

@@ Abstract
The rapid scaling of large vision pretrained models makes fine-tuning tasks more and more difficult on devices with low computational resources. We explore a new visual adaptation paradigm called separated tuning, which treats large pretrained models as standalone feature extractors that run on powerful cloud servers. The fine-tuning carries out on devices which possess only low computational resources (slow CPU, no GPU, small memory, etc.) Existing methods that are potentially suitable for our separated tuning paradigm are discussed. But, three major drawbacks hinder their application in separated tuning: low adaptation capability, large adapter network, and in particular, high information transfer overhead. To address these issues, we propose Minimal Interaction Separated Tuning, or MIST, which reveals that the sum of intermediate features from pretrained models not only has minimal information transfer but also has high adaptation capability. With a lightweight attention-based adaptor network, MIST achieves information transfer efficiency, parameter efficiency, computational and memory efficiency, and at the same time demonstrates competitive results on various visual adaptation benchmarks.
11footnotetext: J. Wu is the corresponding author. This paper was partly supported by the National Natural Science Foundation of China under grant 62276123.
@@ 1 Introduction
Pretraining a large model [10, 9, 3, 22] on general-purpose data and fine-tuning it on specific downstream tasks is emerging as the mainstream methodology for computer vision tasks. Knowledge or representation learned from the pretraining makes adaptation to a target domain much easier than training from scratch.
@@
However, some paradoxes begin to emerge within this process or learning paradigm. One natural issue is: if the downstream dataset is small, we cannot afford to and should not store one new fine-tuned model for each downstream visual task. Parameter-efficient fine-tuning (PEFT) methods [14, 33, 15, 28, 19] are proposed to fine-tune large pretrained models with only a small set of trainable parameters, by either inserting small learnable blocks, adding few learnable tokens, or learning low-rank decomposition of extra parameters. These methods achieve competitive results on various downstream adaptation tasks with less than 1% trainable parameters compared to full fine-tuning.
@@
While PEFT methods partially solved the storage cost of fine-tuned models, other problems remain. As the scale of pretrained models grows with a blazing fast speed, fine-tuning or even inference itself is becoming unaffordable, especially on devices with limited computational resources. On the other hand, the need for fine-tuning and inference on such devices (e.g., personal computers, mobile phones or even cameras) rather than on the cloud AI infrastructure is increasing rapidly.
@@
Figure 1: Illustration of our proposed Separated Tuning. Visual samples for different tasks ($x_{1},x_{2},x_{3},\dots$) are collected and transferred to the cloud server. Within the cloud server, a large pretrained model $B$ acts as a standalone feature extractor, producing intermediate feature sets $(B,x_{i})$. A gather function compresses this set to $\mathrm{Gather}(B,x_{i})$ to achieve minimal interaction but still keeps essential information in $(B,x_{i})$ for downstream task learning. The gathered features for task $i$ is sent to a low-resource device $E_{i}$, where the fine-tuning is performed solely on $E_{i}$ with input $\mathrm{Gather}(B,x_{i})$. During inference, the pretrained model extracts features, gather them to a minimal level, transfers only a small chunk of bytes to the low-resource device, which then makes decisions with small storage, computational, and memory costs.
In this paper, we propose a new visual adaptation paradigm: separated tuning, which enables low-resource devices to fine-tune downstream visual tasks with a small side network (cf. Figure 1). The side network merely receives intermediate features/activations from a pretrained model, such that the huge pretrained network is not required to be (either forward or backward) computed on low-resource devices. Upon the proposed paradigm, we design a new visual adaptation method: Minimal Interaction Separated Tuning, which is based on the side-tuning idea [27, 29]. Methods in this style potentially enables separated tuning, but the interaction (i.e., features transferred) from the pretrained network to the low-resource device is too large for a low-resource device to handle. Our contributions can be summarized as follows:
  • We propose a new visual adaptation paradigm: separated tuning, where the pretrained model works as a service in a cloud server, and low-resource devices fine-tune small side networks to transfer the pretrained large model towards specific downstream tasks.
  • To reduce interaction between server and low-resource devices, we reveal that sum of intermediate features effectively capture essence of the pretrained network, which leads to both minimal interaction and high accuracy.
  • We propose a new visual adaptation method: MIST, as an effective means for separated tuning. MIST is parameter-efficient, GPU memory-efficient, computation efficient, and enjoys low communication overhead between a cloud server and low-resource devices. Experimental results on various datasets show that MIST is comparable with state-of-the art PEFT methods in visual adaptation, but enjoys significantly smaller resource consumptions.
2 New Pipeline: Separated Tuning
@@
As illustrated and described in Figure 1, the idea and structure of separated tuning are pretty simple. Computations are split between a pretrained model $B$ (residing in a cloud server) and fine-tuning modules (which usually happens in edge devices). The pretrained model $B$ is general-purpose and pretrained with abundant data and computing resources, which is too big to infer on a low-resource device. The low-resource device is mostly likely not powerful even to store the model $B$ itself, nor does it has the compute or GPU memory resources to forward compute $B$. The solution is to let $B$ extract features on cloud and transfer them to the low-resource device.
@@
Very important but less observed or discussed, the low-resource device often lacks communication resources, too. That is, in real world it is often difficult or even impossible to transfer large chunks of data between it and the server, e.g., due to bandwidth or energy constraints. Given an input image $x$, $B$ extracts features $(B,x)$ and sends it to the low-resource device. For example, $(B,x)$ can be the set of activations of all intermediate Transformer blocks’ activations, which contains tens of times more bytes to transfer than that in the input $x$ itself. In such a case, communication is too heavy a burden for the low-resource device.
@@
Practical applications, however, are often keen to fine-tune and infer with a model for downstream tasks solely on a low-resource device with limited storage, compute, memory, and communication capabilities.
@@ 2.1 Recipes and Obstacles
Methods that allow to split computations between a large backbone and a small side-network has appeared in the community [27, 29], where the side-network receives features from the backbone but does not send its features back to the computation graph of the backbone. Thus, during fine-tuning, the backbone does not need to compute gradients. This split greatly reduces the compute and GPU memory costs of fine-tuning.
An even simpler recipe is linear evaluation, in which final features of the pretrained model $B$ is coupled with a simple fully connected (FC) layer. Fine-tuning is then learning the FC layer, where the transferred features are tiny, the storage, compute and memory costs are small, too. Partial-$k$ tuning unfreezes the last $k$ layers of $B$ during fine-tuning. When $k$ is small (e.g., $k=1$), it might be able to fit into low-resource devices.
However, these methods do not yet fit the purpose of separated tuning we propose in Figure 1. The set of 5 goals in our separated tuning (small storage/parameters, small compute, small memory footprint, small communication, high accuracy) cannot be met by existing methods. For example,
@@
  • Side-tuning methods [27] produce and transfer large features $(B,x)$, and the side-network may be too big for a low-resource device.
2.2 Minimal Interaction Separated Tuning
Our solution, as depicted in Figure 1, meets these requirements by proposing a gather function $\mathrm{Gather}$, which compresses the information in $(B,x)$ to $\mathrm{Gather}(B,x)$. The gathered output is succinct, hence achieving minimal interaction or communication. So long as $\mathrm{Gather}(B,x)$ distills essential information for the downstream task, the small network on the low-resource device, $E(\mathrm{Gather}(B,x))$ will achieve high accuracy. Because in Figure 1, $E_{i}$ is very small and it does not transfer any features back to $B$, the fine-tuning and inference processes resemble side-tuning methods, which are storage (parameter), compute, and memory efficient, too.
One additional benefit of our pipeline is that the pretrained model is agnostic to downstream tasks. The pretrained model $B$ does not need to store any task-specific parameters or gradients. It simply extract features for any input images. If an input image $x$ is useful in many downstream tasks (correspondingly many low-resource devices), the pretrained model $B$ only needs to compute its features $\mathrm{Gather}(B,x)$ once.
@@ 3 Method
where $z_{0}$ is the embedding of $x$, and $z_{i}$ is the activations/features of the $i$-th Transformer block in $B$, as illustrated in Figure 2. The ‘$\mathrm{Stack}$’ operator stacks or concatenates all tensors $z_{i}$ together. For example, when the pretrained network $B$ is ViT-B, one $3\times 224\times 224$ input image (0.14MB in size) is required to transfer more than 7.5MB to a low-resource device!
Because side tuning methods [27, 29] have achieved state-of-the-art accuracy in visual adaptation tasks, it is a natural hypothesis that $\mathrm{Stack}(z_{0},z_{1},\cdots,z_{N})$ contains enough information for downstream tasks. And, we aim to find a new gather function $\mathrm{Gather^{*}}$ that can most effectively reduce its size but keep the essential information for downstream tasks.
Figure 3: Illustration of MIST and other fine-tuning methods. The Adapter method fine-tunes a model by inserting trainable adapter blocks. LoRA adds trainable low-rank decomposed matrices to parameters. Side-tuning trains a decoupled side network by passing intermediate features to the side network block-wise. Our MIST compresses intermediate features by the gather function $\mathrm{Gather^{*}}$, and passes it to a trainable edge network to adapt to a specific task.
3.1 Gather Function: Additive Feature Integration
@@
Then the input for the $(i+1)$-th side block is $o_{i}+z_{i}$, or equivalently,
@@
To fit into our separated tuning pipeline, $B$ can only send external feature terms to low-resource devices, a requirement only the third term $\Sigma_{l=0}^{i}z_{l}$ satisfies. This term is developed through repeatedly applying the skip connections.
Furthermore, because side network computations ($F_{i}$) are low rank transformation to achieve parameter efficiency [29], a natural conjecture is that this third term (which is determined by the complex model $B$) contains more complicated features compared to the first two terms (which are determined by the simple computations $F_{i}$). In other words, it is reasonable to guess that the third term, $\Sigma_{l=0}^{i}z_{l}$, contains essential information in $in_{i+1}$. Please note that this conjecture is made with respect to any $i$ ($1\leq i\leq N$), not only the final block (i.e., when $i=N$).
@@
This gather function leads to only $1/(N+1)$ communication overhead compared to the one in Eq. 1. Smaller input to low-resource devices also means that the storage, compute, and memory costs will be reduced accordingly. As our experimental results will show, $\mathrm{Gather^{*}}(B,x)$ does contain essential information for downstream tasks, which leads to high accuracy for them.
@@ 3.2 Gather Function: Migrating to Downstream Tasks
One difficulty in handling different downstream tasks are the domain gaps. In some simple tasks with only low-level image patterns, high-level features extracted by a pretrained model may be harmful for adaptation. In such cases, gathering all the intermediate features altogether is suboptimal.
So we migrate the extracted features to specific downstream tasks by only gathering features from the first $k$ blocks, which leads to the final form of our minimal interaction gather function:
@@
This final form introduces a hyperparameter, $k$, which needs to be set before separated tuning. We select the best $k$ from a candidate set by using the validation set. The candidate set is formed by $N$, $43N/$ 4 and $N/2$. For example, for ViT-B with 12 layers, we pick $k$ from the set $\{6,9,12\}$.
@@
MIST (ours) 0.577 0.38 17.57 00.09 0.57 0.45 81.4 86.4 62.4 76.7
Table 1: Results on VTAB-1K. “Params”: number of trainable parameters. “MACs Cloud”, “MACs Edge”: multiply–accumulate operations for cloud and low-resource devices during fine-tuning. “GPU Mem Cloud”, “GPU Mem Edge”: GPU memory footprint within cloud and low-resource devices, respectively during fine-tuning with batch size 32. “Nat.”, “Spe.”, “Str.”: average accuracy for “Natural”, “Specialized”, “Structured” task groups. “Avg.”: group-wise average accuracy. Best results are in boldface.
3.3 The Proposed MIST Method
With $\mathrm{Gather^{*}}$, our Minimal Interaction Separated Tuning (MIST) method becomes clear and easy to describe. It contains a pretrained base vision transformer network $B$ and a low-rank attention edge network $\mathrm{LAE}$ (which will be introduced latter).
$Z_{\mathrm{mix}}=\mathrm{Gather^{*}}(B,x,k)\,.$ (7)
$Z_{\mathrm{out}}=\mathrm{LAE}(Z_{\mathrm{mix}})\,.$ (8)
From the result in [29], attention block with low-dimensional QKV projection is an effective and efficient structure for visual adaptation. Our experimental results also show that it is very effective for separated tuning. So we adopt this structure for edge network.
As shown in Figure 3, $\mathrm{LAE}$ contains $L$ blocks. Each block contains a layer norm [1] and a low-rank attention, where low-rank attention refers to self-attention with an extremely low QKV hidden dimension $r$ (e.g., hidden dimension equals 16 for feature dimension equals 768). Skip connection is also applied in each block. In practice, we take $L=4$ and $r=32$ for all downstream tasks, and find that such a small edge network is already enough for most visual adaptation tasks.
4 Experiments